# Mixture-of-Subspaces in Low-Rank Adaptation

### v2

@@ 1 Introduction
Figure 1: Comparison between vanilla LoRA and proposed MoSLoRA. In MoSLoRA, we employ learnable weights to mix more subspaces with negligible parameters (i.e. $(d_{1}+d_{2}+r)r$ vs $(d_{1}+d_{2})r$ and $d_{1}+d_{2}\gg r$ typically).
@@
Motivated by the observation, we further revisit the two-subspaces-mixing strategy in a more fine-grained (rank=1) view and composed view. In short, such a strategy equals inserting a mixer matrix between $\mathbf{A}$ and $\mathbf{B}$, which is a fixed butterfly factor (Dao et al., 2019). Meanwhile, vanilla LoRA can be considered as a special case with a fixed identity matrix being the mixer. Therefore, we propose MoSLoRA, a simple yet effective method, which employs a learnable mixer to fuse more subspaces and more flexibly. As shown in Figure 1, we adapt the mixer $\mathbf{W}$ to fuse all the possible subspaces (i.e. $A_{i}B_{j}$). Compared to LoRA, MoSLoRA requires negligible extra parameters since $d_{1}+d_{2}\gg r$. Similarly to LoRA, MoSLoRA can also be merged into the original weights, and thus introduce no latency during inference.
@@ 2.1 LoRA and Subspace View
In this paper, we decompose LoRA into subspaces via structural re-parameterization, where the subspaces are defined as parallel components with smaller rank values. Figure 2 part a shows the procedure for two subspaces. Specifically, we decompose the $\mathbf{A}$ into two parts (i.e. $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$) by column, and $\mathbf{B}$ by row to get $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$. Therefore, we can easily get that:
@@ 2.2 Mixing Two Subspaces
$=$ $x(\mathbf{A}_{1}\mathbf{B}_{1}+\mathbf{A}_{2}\mathbf{B}_{2}+{\color[rgb]{1,0,0}\mathbf{A}_{1}\mathbf{B}_{2}}+{\color[rgb]{1,0,0}\mathbf{A}_{2}\mathbf{B}_{1}}).$
@@ 3.1 More Fine-grained Subspace
Motivated by the observation that mixing two subspaces would lead to better performance, we revisit the two-subspaces-mixing LoRA in view of more fine-grained subspace (i.e. rank=1). Specifically, we decompose the $\mathbf{A}\in\mathbb{R}^{d_{1}\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times d_{2}}$ into $r$ subspaces (rank=1), which can be formulated as:
@@
$\mathbf{A}{\color[rgb]{1,0,0}\mathbf{W}}\mathbf{B},$ (7)
@@ 3.2 Initialization Strategies for Mixer
In the proposed MoSLoRA, we employ a trainable mixer to fuse all possible subspaces. However, the system of MoSLoRA is linear, and thus a bad initialization can still hamper the learning (He et al., 2015). In MoSLoRA, we follow the setting in LoRA and initialize $\mathbf{A}$ using a Kaiming uniform distribution  [^2] In the code of LoRA, they use Kaiming uniform initialization rather than Gaussian distribution claimed in the paper. and $\mathbf{B}$ as a zero matrix. For the mixer weight $\mathbf{W}$, we compare various initialization strategies, including zero matrix, identity matrix, normal distribution, orthogonal matrix (Saxe et al., 2014), and Kaiming uniform distribution (He et al., 2015). Hyperparameters for finetuning can be found at Appendix B.1.
@@ 3.3 Relation with Mixture-of-Experts
Mixture-of-Experts (MoE) methods aim to partition a set of parameters into experts and route input samples to specific experts during training and inference (Fedus et al., 2022a). Typically, they employ a router to generate scores for each expert based on the input, and then select top-k experts (Fedus et al., 2022b; Lepikhin et al., 2021; DeepSeek-AI, 2024). In this paper, we propose MoSLoRA to mix the subspaces in LoRA, where the $w_{ij}$ in the mixer can be considered as the weight to compose subspace $\mathbf{A}_{i}\mathbf{B}_{j}$. However, the differences between MoSLoRA and MoE methods are as follows:
@@ 4.3 Subject-driven Generation
Figure 6 shows one case of a dog from the DreamBooth dataset (Ruiz et al., 2023). We finetune the SDXL555stable-diffusion-xl-base-1.0 model (Podell et al., 2023) via LoRA and MoSLoRA. In MoSLoRA, the mixer is initialized as an orthogonal matrix. During finetuning, the learning rate is 1e-4, and the batch size is 4. We train the model for 500 steps, which costs around 16 minutes using 1 80G A100 GPU. During generation, we infer 50 steps for the given prompts. Compared to vanilla LoRA, we can find that our proposed MoSLoRA captures more details of the subject and better conforms to the given prompt. Specifically, MoSLoRA learns more details about the given cat, including the color of the hairs around the neck and the shape of the paw. Meanwhile, the images from MoSLoRA are more consistent with the given prompts, such as the wet (thus clumped) hair and the floating gesture (spread hands).
@@
We also perform human evaluation on the generated images. First, we choose four subjects (i.e. cat, dog, grey sloth plushie, and can) from the DreamBooth dataset (Ruiz et al., 2023) and fine-tune the SDXL model, respectively. Then, we randomly select 8 prompts to generate the corresponding images. After that, 15 human experts are asked to independently score win/tie/loss for the paired images from LoRA and MoSLoRA. During evaluation, we shuffle these pairs and keep that these experts do not know the source model of each image. We employ two metrics, including 1) subject similarity defined as the similarity between subjects from generated images and given images, and 2) prompt consistency defined as the consistency among prompts and generated images. Table 6 reports the average score for all the images. We can find that MoSLoRA outperforms LoRA on both metrics. In particular, MoSLoRA gets an average winning ratio of 45.1% on prompt consistency, which is 34.3% than LoRA. Please refer to Appendix D for the detailed prompts and corresponding generated images from LoRA and MoSLoRA.
@@ 6 Conclusion
This work proposes a novel MoSLoRA method for parameter-efficient fine-tuning. We first decompose the LoRA into subspaces and find that simply mixing the half-rank subspaces would lead to better performance. After that, we revisit vanilla LoRA and two-subspaces-mixing strategy in a more fine-grained view (i.e. rank=1), thus unifying both methods as employing an extra fixed mixer. Therefore, we propose MoSLoRA, which employs a learnable mixer to fuse more information and more flexibly. The mixer requires negligible extra parameters and computing costs. Experimental results on commonsense reasoning tasks, visual instruction tuning tasks, and subject-driven generation tasks demonstrate the effectiveness and robustness of the proposed MoSLoRA. For future work, we would consider applying MoSLoRA for more tasks. Finding a task-specific way to initialize the mixer for faster convergence would be another interesting topic.

### v3

@@ 1 Introduction
Figure 1: Comparison between vanilla LoRA and proposed MoSLoRA. In MoSLoRA, we employ learnable weights to mix more subspaces with negligible parameters (i.e., $(d_{1}+d_{2}+r)r$ vs $(d_{1}+d_{2})r$ and $d_{1}+d_{2}\gg r$ typically).
@@
Motivated by the observation, we further revisit the two-subspaces-mixing strategy in a more fine-grained (rank=1) view and composed view. In short, such a strategy equals inserting a mixer matrix between $\mathbf{A}$ and $\mathbf{B}$, which is a fixed butterfly factor (Dao et al., 2019). Meanwhile, vanilla LoRA can be considered as a special case with a fixed identity matrix being the mixer. Therefore, we propose MoSLoRA, a simple yet effective method, which employs a learnable mixer to fuse more subspaces and more flexibly. As shown in Figure 1, we adapt the mixer $\mathbf{W}$ to fuse all the possible subspaces (i.e., $A_{i}B_{j}$). Compared to LoRA, MoSLoRA requires negligible extra parameters since $d_{1}+d_{2}\gg r$. Similarly to LoRA, MoSLoRA can also be merged into the original weights, and thus introduce no latency during inference.
@@ 2.1 LoRA and Subspace View
In this paper, we decompose LoRA into subspaces via structural re-parameterization, where the subspaces are defined as parallel components with smaller rank values. Figure 2 part a shows the procedure for two subspaces. Specifically, we decompose the $\mathbf{A}$ into two parts (i.e., $\mathbf{A}_{1}$ and $\mathbf{A}_{2}$) by column, and $\mathbf{B}$ by row to get $\mathbf{B}_{1}$ and $\mathbf{B}_{2}$. Therefore, we can easily get that:
@@ 2.2 Mixing Two Subspaces
$=$ $x(\mathbf{A}_{1}\mathbf{B}_{1}+\mathbf{A}_{2}\mathbf{B}_{2}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{A}_{1}\mathbf{B}_{2}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{A}_{2}\mathbf{B}_{1}}).$
@@ 3.1 More Fine-grained Subspace
Motivated by the observation that mixing two subspaces would lead to better performance, we revisit the two-subspaces-mixing LoRA in view of more fine-grained subspace (i.e., rank=1). Specifically, we decompose the $\mathbf{A}\in\mathbb{R}^{d_{1}\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times d_{2}}$ into $r$ subspaces (rank=1), which can be formulated as:
@@
$\mathbf{A}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{W}}\mathbf{B},$ (7)
@@ 3.2 Initialization Strategies for Mixer
In the proposed MoSLoRA, we employ a trainable mixer to fuse all possible subspaces. However, the system of MoSLoRA is linear, and a bad initialization hampers the learning (He et al., 2015). In MoSLoRA, we follow the setting in LoRA and initialize $\mathbf{A}$ using a Kaiming uniform distribution  [^2] In the code of LoRA, they use Kaiming uniform initialization rather than Gaussian distribution claimed in the paper. and $\mathbf{B}$ as a zero matrix. For the mixer weight $\mathbf{W}$, we compare various initialization strategies, including zero matrix, identity matrix, normal distribution, orthogonal matrix (Saxe et al., 2014), and Kaiming uniform distribution (He et al., 2015). Hyperparameters for finetuning can be found at Appendix B.1.
@@ 3.3 Relation with Mixture-of-Experts
Mixture-of-Experts (MoE) methods aim to partition a set of parameters into experts and route input samples to specific experts during training and inference (Fedus et al., 2022a; Shi et al., 2024). Typically, they employ a router to generate scores for each expert based on the input, and then select top-k experts (Fedus et al., 2022b; Lepikhin et al., 2021; DeepSeek-AI, 2024). In this paper, we propose MoSLoRA to mix the subspaces in LoRA, where the $w_{ij}$ in the mixer can be considered as the weight to compose subspace $\mathbf{A}_{i}\mathbf{B}_{j}$. However, the differences between MoSLoRA and MoE methods are as follows:
@@ 4.3 Subject-driven Generation
Figure 6 shows one case of a cat from the DreamBooth dataset (Ruiz et al., 2023). We finetune the SDXL555stable-diffusion-xl-base-1.0 model (Podell et al., 2023) via LoRA and MoSLoRA. In MoSLoRA, the mixer is initialized as an orthogonal matrix. During finetuning, the learning rate is 1e-4, and the batch size is 4. We train the model for 500 steps, which costs around 16 minutes using 1 80G A100 GPU. During generation, we infer 50 steps for the given prompts. Compared to vanilla LoRA, we can find that our proposed MoSLoRA captures more details of the subject and better conforms to the given prompt. Specifically, MoSLoRA learns more details about the given cat, including the color of the hairs around the neck and the shape of the paw. Meanwhile, the images from MoSLoRA are more consistent with the given prompts, such as the wet (thus clumped) hair and the floating gesture (spread hands).
@@
We also perform human evaluation on the generated images. First, we choose four subjects (i.e., cat, dog, grey sloth plushie, and can) from the DreamBooth dataset (Ruiz et al., 2023) and fine-tune the SDXL model, respectively. Then, we randomly select 8 prompts to generate the corresponding images. After that, 15 human experts are asked to independently score win/tie/loss for the paired images from LoRA and MoSLoRA. During evaluation, we shuffle these pairs and keep that these experts do not know the source model of each image. We employ two metrics, including 1) subject similarity defined as the similarity between subjects from generated images and given images, and 2) prompt consistency defined as the consistency among prompts and generated images. Table 6 reports the average score for all the images. We can find that MoSLoRA outperforms LoRA on both metrics. In particular, MoSLoRA gets an average winning ratio of 45.1% on prompt consistency, which is 34.3% than LoRA. Please refer to Appendix D for the detailed prompts and corresponding generated images from LoRA and MoSLoRA.
@@ 6 Conclusion
This work proposes a novel MoSLoRA method for parameter-efficient fine-tuning. We first decompose the LoRA into subspaces and find that simply mixing the half-rank subspaces would lead to better performance. After that, we revisit vanilla LoRA and two-subspaces-mixing strategy in a more fine-grained view (i.e., rank=1), thus unifying both methods as employing an extra fixed mixer. Therefore, we propose MoSLoRA, which employs a learnable mixer to fuse more information and more flexibly. The mixer requires negligible extra parameters and computing costs. Experimental results on commonsense reasoning tasks, visual instruction tuning tasks, and subject-driven generation tasks demonstrate the effectiveness and robustness of the proposed MoSLoRA. For future work, we would consider applying MoSLoRA for more tasks. Finding a task-specific way to initialize the mixer for faster convergence would be another interesting topic.
@@
Acknowledgements
We thank all anonymous reviewers for their constructive feedback on improving our paper. We thank Zenan Xu, Chaofan Tao, Chenchen Ding, Shuqi Wang, and Zhengwu Liu for their fruitful discussions. This work was supported by the Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC), Hong Kong SAR.
@@ Appendix D Cases of Generated Images
Appendix E Analysis of $\mathbf{W}$ in MoSLoRA
E.1 Vanilla MoSLoRA
For an arbitrary input $x$, we have:
$y$ $=x\mathbf{W}_{merge};\mathbf{W}_{merge}=\mathbf{W}_{0}+\mathbf{A}\mathbf{W}\mathbf{B},$ (10)
where the $\mathbf{W}_{0}$ is frozen during training. Then we have:
$\frac{\partial y}{\partial\mathbf{A}}=\frac{\partial y}{\partial\mathbf{W}_{merge}}\mathbf{B}^{T}\mathbf{W}^{T};\frac{\partial y}{\partial\mathbf{W}}$ $=\mathbf{A}^{T}\frac{\partial y}{\partial\mathbf{W}_{merge}}\mathbf{B}^{T};\frac{\partial y}{\partial\mathbf{B}}=\mathbf{W}^{T}\mathbf{A}^{T}\frac{\partial y}{\partial\mathbf{W}_{merge}}$ (11)
Denote the learning rate as $\eta$, the updating process is:
$\mathbf{A}\leftarrow\mathbf{A}-\eta\frac{\partial y}{\partial\mathbf{A}}=\mathbf{A}-\eta\frac{\partial y}{\partial\mathbf{W}_{merge}}\mathbf{B}^{T}\mathbf{W}^{T}$ (12)
The process is similar for $\mathbf{W}$ and $\mathbf{B}$. Let $\Delta=\frac{\partial y}{\partial\mathbf{W}_{merge}}$. Thus, the weight of the updated LoRA branch would be:
$\mathbf{W}_{LoRA}$ $=(\mathbf{A}-\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T})(\mathbf{W}-\eta\mathbf{A}^{T}\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$ (13)
$=(\mathbf{A}\mathbf{W}-\eta\mathbf{A}\mathbf{A}^{T}\Delta\mathbf{B}^{T}-\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T}\mathbf{W}+\eta^{2}\Delta\mathbf{B}^{T}\mathbf{W}^{T}\mathbf{A}^{T}\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$
E.2 Merge $\mathbf{A}$ and $\mathbf{W}$
Denote $\hat{\mathbf{A}}=\mathbf{A}\mathbf{W}$. It means that we initialize $\hat{\mathbf{A}}$ as the same as $\mathbf{A}\mathbf{W}$. The output is the same:
$y$ $=x\mathbf{W}_{merge};\mathbf{W}_{merge}=\mathbf{W}_{0}+\hat{\mathbf{A}}\mathbf{B}=\mathbf{W}_{0}+\mathbf{A}\mathbf{W}\mathbf{B}.$ (14)
However, the corresponding gradients would be:
$\frac{\partial y}{\partial\hat{\mathbf{A}}}=\frac{\partial y}{\partial\mathbf{W}_{merge}}\mathbf{B}^{T}=\Delta\mathbf{B}^{T};\frac{\partial y}{\partial\mathbf{B}}=\hat{\mathbf{A}}^{T}\frac{\partial y}{\partial\mathbf{W}_{merge}}=\hat{\mathbf{A}}^{T}\Delta$ (15)
Based on that, we can get the updated LoRA after updating the parameters:
$\hat{\mathbf{W}}_{LoRA}$ $=(\hat{\mathbf{A}}-\eta\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\hat{\mathbf{A}}^{T}\Delta)$ (16)
$=(\mathbf{A}\mathbf{W}-\eta\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$
E.3 Comparison
Comparing Equation 13 and 16, we can conclude that the updated weights are not the same, since
$\hat{\mathbf{W}}_{LoRA}-\mathbf{W}_{LoRA}$ $=(-\eta\Delta\mathbf{B}^{T}+\eta\mathbf{A}\mathbf{A}^{T}\Delta\mathbf{B}^{T}+\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T}\mathbf{W}-\eta^{2}\Delta\mathbf{B}^{T}\mathbf{W}^{T}\mathbf{A}^{T}\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$ (17)
$=(\eta(\mathbf{A}-\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T})\mathbf{A}^{T}\Delta\mathbf{B}^{T}+\eta\Delta\mathbf{B}^{T}(\mathbf{W}^{T}\mathbf{W}-\mathbf{I}))(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)\neq\mathbf{0}.$
E.4 Fix $\mathbf{W}$ as Orthogonal Matrix
If we fix $\mathbf{W}$ as orthogonal matrix and do not update (i.e., $\mathbf{W}\mathbf{W}^{T}=\mathbf{I}$), the updated LoRA would be:
$\mathbf{W}^{\mathbf{I}}_{LoRA}$ $=(\mathbf{A}-\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T})\mathbf{W}(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$ (18)
$=(\mathbf{A}\mathbf{W}-\eta\Delta\mathbf{B}^{T}\mathbf{W}^{T}\mathbf{W})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)$
$=(\mathbf{A}\mathbf{W}-\eta\Delta\mathbf{B}^{T})(\mathbf{B}-\eta\mathbf{W}^{T}\mathbf{A}^{T}\Delta)=\hat{\mathbf{W}}_{LoRA}$
E.5 Conclusion
Though mathematically equivalent initialized, the optimization process would be different if $\mathbf{W}$ is learnable. Specifically, the optimization process would be the same i.i.f $\mathbf{W}$ is a fixed orthogonal matrix.